OpenStudy (anonymous):

Eliminate the parameter. x = t2 + 1, y = t2 - 1

OpenStudy (amoodarya):

x-y=2

OpenStudy (anonymous):

\[ x = t^2 + 1 \quad \quad\quad y = t^2 - 1 \\ t^2 = x - 1 \\ y = t^2 - 1 \quad \to \quad y = (x - 1) - 1 \\ y = x - 2 \]

OpenStudy (anonymous):

y = x + 2, x ≥ 1 y = x - 2, x ≥ 1 y = x2 - 2, x ≥ 1 y = x2 + 2, x ≥ 1

OpenStudy (anonymous):

so its b

OpenStudy (anonymous):

yes

OpenStudy (amoodarya):

x=t^2+1>=1

OpenStudy (anonymous):

Eliminate the parameter. x = square root of t, y = 4t + 1

OpenStudy (anonymous):

y = 4square root of x - 1, x ≥ 0 ?

OpenStudy (amoodarya):

x^2=t so put t in y=4t+1

OpenStudy (anonymous):

y = 4square root of x + 1, x ≥ 0 y = 4square root of x - 1, x ≥ 0 y = 4x2 - 1 y = 4x2 + 1

OpenStudy (amoodarya):

y=4(x^2)-1

OpenStudy (anonymous):

Eliminate the parameter. x = 3t, y = t + 7

OpenStudy (amoodarya):

t=y-7 put it in x=3t y=3(y-7) then simplify